\begin{thebibliography}{69}
\providecommand{\natexlab}[1]{#1}
\providecommand{\url}[1]{\texttt{#1}}
\expandafter\ifx\csname urlstyle\endcsname\relax
  \providecommand{\doi}[1]{doi: #1}\else
  \providecommand{\doi}{doi: \begingroup \urlstyle{rm}\Url}\fi

\bibitem[Ackerman et~al.(2021)Ackerman, Raz, Zalmanovici, and
  Zlotnick]{ackerman2021automatically}
S.~Ackerman, O.~Raz, M.~Zalmanovici, and A.~Zlotnick.
\newblock Automatically detecting data drift in machine learning classifiers.
\newblock \emph{arXiv preprint arXiv:2111.05672}, 2021.

\bibitem[Adams and MacKay(2007)]{adams2007bayesian}
R.~P. Adams and D.~J. MacKay.
\newblock Bayesian online changepoint detection.
\newblock \emph{arXiv preprint arXiv:0710.3742}, 2007.

\bibitem[Baena-Garc{\i}a et~al.(2006)Baena-Garc{\i}a, del Campo-{\'A}vila,
  Fidalgo, Bifet, Gavalda, and Morales-Bueno]{baena2006early}
M.~Baena-Garc{\i}a, J.~del Campo-{\'A}vila, R.~Fidalgo, A.~Bifet, R.~Gavalda,
  and R.~Morales-Bueno.
\newblock Early drift detection method.
\newblock In \emph{Fourth international workshop on knowledge discovery from
  data streams}, volume~6, pages 77--86, 2006.

\bibitem[Barros et~al.(2017)Barros, Cabral, Gon{\c{c}}alves~Jr, and
  Santos]{barros2017rddm}
R.~S. Barros, D.~R. Cabral, P.~M. Gon{\c{c}}alves~Jr, and S.~G. Santos.
\newblock Rddm: Reactive drift detection method.
\newblock \emph{Expert Systems with Applications}, 90:\penalty0 344--355, 2017.

\bibitem[Bayram et~al.(2022)Bayram, Ahmed, and Kassler]{bayram2022concept}
F.~Bayram, B.~S. Ahmed, and A.~Kassler.
\newblock From concept drift to model degradation: An overview on
  performance-aware drift detectors.
\newblock \emph{Knowledge-Based Systems}, page 108632, 2022.

\bibitem[Behnel et~al.(2010)Behnel, Bradshaw, Citro, Dalcin, Seljebotn, and
  Smith]{behnel2010cython}
S.~Behnel, R.~Bradshaw, C.~Citro, L.~Dalcin, D.~S. Seljebotn, and K.~Smith.
\newblock Cython: The best of both worlds.
\newblock \emph{Computing in Science \& Engineering}, 13\penalty0 (2):\penalty0
  31--39, 2010.

\bibitem[Bhattacharyya(1946)]{bhattacharyya1946measure}
A.~Bhattacharyya.
\newblock On a measure of divergence between two multinomial populations.
\newblock \emph{Sankhy{\=a}: the indian journal of statistics}, pages 401--406,
  1946.

\bibitem[Bifet and Gavalda(2007)]{bifet2007learning}
A.~Bifet and R.~Gavalda.
\newblock Learning from time-changing data with adaptive windowing.
\newblock In \emph{Proceedings of the 2007 SIAM international conference on
  data mining}, pages 443--448. SIAM, 2007.

\bibitem[Bifet et~al.(2010)Bifet, Holmes, Pfahringer, Kranen, Kremer, Jansen,
  and Seidl]{bifet2010moa}
A.~Bifet, G.~Holmes, B.~Pfahringer, P.~Kranen, H.~Kremer, T.~Jansen, and
  T.~Seidl.
\newblock Moa: Massive online analysis, a framework for stream classification
  and clustering.
\newblock In \emph{Proceedings of the first workshop on applications of pattern
  analysis}, pages 44--50. PMLR, 2010.

\bibitem[Bressler et~al.(2022)Bressler, Gabbay, Chorev, Tannor, Perlmutter,
  Hutnik, Liberman, Ben~Israel, and Romanyshyn]{deepchecks}
N.~Bressler, I.~Gabbay, S.~Chorev, P.~Tannor, M.~Perlmutter, N.~Hutnik,
  J.~Liberman, D.~Ben~Israel, and Y.~Romanyshyn.
\newblock Deepchecks: Comprehensive testing and validation of ml models and
  data, 2022.
\newblock URL \url{https://www.deepchecks.com/}.
\newblock Software available from https://github.com/deepchecks/deepchecks.

\bibitem[Bu et~al.(2016)Bu, Alippi, and Zhao]{bu2016pdf}
L.~Bu, C.~Alippi, and D.~Zhao.
\newblock A pdf-free change detection test based on density difference
  estimation.
\newblock \emph{IEEE transactions on neural networks and learning systems},
  29\penalty0 (2):\penalty0 324--334, 2016.

\bibitem[Chollet et~al.(2015)]{chollet2015keras}
F.~Chollet et~al.
\newblock Keras.
\newblock \url{https://keras.io}, 2015.

\bibitem[Cobb and Van~Looveren(2022)]{cobb2022context}
O.~Cobb and A.~Van~Looveren.
\newblock Context-aware drift detection.
\newblock In \emph{International Conference on Machine Learning}, pages
  4087--4111. PMLR, 2022.

\bibitem[Cram{\'e}r(1928)]{cramer1928composition}
H.~Cram{\'e}r.
\newblock On the composition of elementary errors: First paper: Mathematical
  deductions.
\newblock \emph{Scandinavian Actuarial Journal}, 1928\penalty0 (1):\penalty0
  13--74, 1928.

\bibitem[Dasu et~al.(2006)Dasu, Krishnan, Venkatasubramanian, and
  Yi]{dasu2006information}
T.~Dasu, S.~Krishnan, S.~Venkatasubramanian, and K.~Yi.
\newblock An information-theoretic approach to detecting changes in
  multi-dimensional data streams.
\newblock In \emph{In Proc. Symp. on the Interface of Statistics, Computing
  Science, and Applications}. Citeseer, 2006.

\bibitem[Ditzler and Polikar(2011)]{ditzler2011hellinger}
G.~Ditzler and R.~Polikar.
\newblock Hellinger distance based drift detection for nonstationary
  environments.
\newblock In \emph{2011 IEEE symposium on computational intelligence in dynamic
  and uncertain environments (CIDUE)}, pages 41--48. IEEE, 2011.

\bibitem[dos Reis et~al.(2016)dos Reis, Flach, Matwin, and
  Batista]{dos2016fast}
D.~M. dos Reis, P.~Flach, S.~Matwin, and G.~Batista.
\newblock Fast unsupervised online drift detection using incremental
  kolmogorov-smirnov test.
\newblock In \emph{Proceedings of the 22nd ACM SIGKDD international conference
  on knowledge discovery and data mining}, pages 1545--1554, 2016.

\bibitem[Dral(2020)]{evidently}
E.~Dral.
\newblock An open-source framework to evaluate, test and monitor ml models in
  production.
\newblock \url{https://github.com/evidentlyai/evidently}, 2020.
\newblock {Online; accessed 23-7-2023}.

\bibitem[Falcon and {The PyTorch Lightning
  team}(2019)]{Falcon_PyTorch_Lightning_2019}
W.~Falcon and {The PyTorch Lightning team}.
\newblock {PyTorch Lightning}, Mar. 2019.
\newblock URL \url{https://github.com/Lightning-AI/lightning}.

\bibitem[Frias-Blanco et~al.(2014)Frias-Blanco, del Campo-{\'A}vila,
  Ramos-Jimenez, Morales-Bueno, Ortiz-Diaz, and
  Caballero-Mota]{frias2014online}
I.~Frias-Blanco, J.~del Campo-{\'A}vila, G.~Ramos-Jimenez, R.~Morales-Bueno,
  A.~Ortiz-Diaz, and Y.~Caballero-Mota.
\newblock Online and non-parametric drift detection methods based on
  hoeffding’s bounds.
\newblock \emph{IEEE Transactions on Knowledge and Data Engineering},
  27\penalty0 (3):\penalty0 810--823, 2014.

\bibitem[Gama et~al.(2004)Gama, Medas, Castillo, and
  Rodrigues]{gama2004learning}
J.~Gama, P.~Medas, G.~Castillo, and P.~Rodrigues.
\newblock Learning with drift detection.
\newblock In \emph{Brazilian symposium on artificial intelligence}, pages
  286--295. Springer, 2004.

\bibitem[Gama et~al.(2013)Gama, Sebastiao, and Rodrigues]{gama2013evaluating}
J.~Gama, R.~Sebastiao, and P.~P. Rodrigues.
\newblock On evaluating stream learning algorithms.
\newblock \emph{Machine learning}, 90\penalty0 (3):\penalty0 317--346, 2013.

\bibitem[Gama et~al.(2014)Gama, {\v{Z}}liobait{\.e}, Bifet, Pechenizkiy, and
  Bouchachia]{gama2014survey}
J.~Gama, I.~{\v{Z}}liobait{\.e}, A.~Bifet, M.~Pechenizkiy, and A.~Bouchachia.
\newblock A survey on concept drift adaptation.
\newblock \emph{ACM computing surveys (CSUR)}, 46\penalty0 (4):\penalty0 1--37,
  2014.

\bibitem[Goldenberg and Webb(2019)]{goldenberg2019survey}
I.~Goldenberg and G.~I. Webb.
\newblock Survey of distance measures for quantifying concept drift and shift
  in numeric data.
\newblock \emph{Knowledge and Information Systems}, 60\penalty0 (2):\penalty0
  591--615, 2019.

\bibitem[Gretton et~al.(2012)Gretton, Borgwardt, Rasch, Sch{{\"o}}lkopf, and
  Smola]{JMLR:v13:gretton12a}
A.~Gretton, K.~M. Borgwardt, M.~J. Rasch, B.~Sch{{\"o}}lkopf, and A.~Smola.
\newblock A kernel two-sample test.
\newblock \emph{Journal of Machine Learning Research}, 13\penalty0
  (25):\penalty0 723--773, 2012.
\newblock URL \url{http://jmlr.org/papers/v13/gretton12a.html}.

\bibitem[Gupta(2023)]{uptrain}
V.~Gupta.
\newblock An open-source framework to observe ml applications, built for
  engineers.
\newblock \url{https://github.com/uptrain-ai/uptrain}, 2023.
\newblock {Online; accessed 23-7-2023}.

\bibitem[Halford et~al.(2019)Halford, Bolmier, Sourty, Vaysse, and
  Zouitine]{creme}
M.~Halford, G.~Bolmier, R.~Sourty, R.~Vaysse, and A.~Zouitine.
\newblock {creme}, a {P}ython library for online machine learning, 2019.
\newblock URL \url{https://github.com/MaxHalford/creme}.

\bibitem[Hellinger(1909)]{hellinger1909neue}
E.~Hellinger.
\newblock Neue begr{\"u}ndung der theorie quadratischer formen von
  unendlichvielen ver{\"a}nderlichen.
\newblock \emph{Journal f{\"u}r die reine und angewandte Mathematik},
  1909\penalty0 (136):\penalty0 210--271, 1909.

\bibitem[Huang et~al.(2014)Huang, Koh, Dobbie, and Pears]{huang2014detecting}
D.~T.~J. Huang, Y.~S. Koh, G.~Dobbie, and R.~Pears.
\newblock Detecting volatility shift in data streams.
\newblock In \emph{2014 IEEE International Conference on Data Mining}, pages
  863--868. IEEE, 2014.

\bibitem[Khamassi et~al.(2018)Khamassi, Sayed-Mouchaweh, Hammami, and
  Gh{\'e}dira]{khamassi2018discussion}
I.~Khamassi, M.~Sayed-Mouchaweh, M.~Hammami, and K.~Gh{\'e}dira.
\newblock Discussion and review on evolving data streams and concept drift
  adapting.
\newblock \emph{Evolving systems}, 9\penalty0 (1):\penalty0 1--23, 2018.

\bibitem[Kullback and Leibler(1951)]{kullback1951information}
S.~Kullback and R.~A. Leibler.
\newblock On information and sufficiency.
\newblock \emph{The annals of mathematical statistics}, 22\penalty0
  (1):\penalty0 79--86, 1951.

\bibitem[Leigh et~al.(2022)Leigh, Schill, Lindsay, Srivastava, McNamara, and
  Jarmale]{menelaus}
N.~Leigh, T.~Schill, I.~Lindsay, A.~Srivastava, K.~P. McNamara, and S.~Jarmale.
\newblock Menelaus.
\newblock \url{https://github.com/mitre/menelaus}, 2022.
\newblock {Online; accessed 23-7-2023}.

\bibitem[Lin(1991)]{lin1991divergence}
J.~Lin.
\newblock Divergence measures based on the shannon entropy.
\newblock \emph{IEEE Transactions on Information theory}, 37\penalty0
  (1):\penalty0 145--151, 1991.

\bibitem[Lindstrom et~al.(2013)Lindstrom, Mac~Namee, and
  Delany]{lindstrom2013drift}
P.~Lindstrom, B.~Mac~Namee, and S.~J. Delany.
\newblock Drift detection using uncertainty distribution divergence.
\newblock \emph{Evolving Systems}, 4\penalty0 (1):\penalty0 13--25, 2013.

\bibitem[Liu et~al.(2018)Liu, Lu, Liu, and Zhang]{liu2018accumulating}
A.~Liu, J.~Lu, F.~Liu, and G.~Zhang.
\newblock Accumulating regional density dissimilarity for concept drift
  detection in data streams.
\newblock \emph{Pattern Recognition}, 76:\penalty0 256--272, 2018.

\bibitem[Liu et~al.(2020)Liu, Xu, Lu, Zhang, Gretton, and
  Sutherland]{liu2020learning}
F.~Liu, W.~Xu, J.~Lu, G.~Zhang, A.~Gretton, and D.~J. Sutherland.
\newblock Learning deep kernels for non-parametric two-sample tests.
\newblock In \emph{International conference on machine learning}, pages
  6316--6326. PMLR, 2020.

\bibitem[Lopez-Paz and Oquab(2016)]{lopez2016revisiting}
D.~Lopez-Paz and M.~Oquab.
\newblock Revisiting classifier two-sample tests.
\newblock \emph{arXiv preprint arXiv:1610.06545}, 2016.

\bibitem[Mann and Whitney(1947)]{mann1947test}
H.~B. Mann and D.~R. Whitney.
\newblock On a test of whether one of two random variables is stochastically
  larger than the other.
\newblock \emph{The annals of mathematical statistics}, pages 50--60, 1947.

\bibitem[Massey~Jr(1951)]{massey1951kolmogorov}
F.~J. Massey~Jr.
\newblock The kolmogorov-smirnov test for goodness of fit.
\newblock \emph{Journal of the American statistical Association}, 46\penalty0
  (253):\penalty0 68--78, 1951.

\bibitem[Montiel et~al.(2018)Montiel, Read, Bifet, and
  Abdessalem]{JMLR:v19:18-251}
J.~Montiel, J.~Read, A.~Bifet, and T.~Abdessalem.
\newblock Scikit-multiflow: A multi-output streaming framework.
\newblock \emph{Journal of Machine Learning Research}, 19\penalty0
  (72):\penalty0 1--5, 2018.
\newblock URL \url{http://jmlr.org/papers/v19/18-251.html}.

\bibitem[Montiel et~al.(2021)Montiel, Halford, Mastelini, Bolmier, Sourty,
  Vaysse, Zouitine, Gomes, Read, Abdessalem, and Bifet]{JMLR:v22:20-1380}
J.~Montiel, M.~Halford, S.~M. Mastelini, G.~Bolmier, R.~Sourty, R.~Vaysse,
  A.~Zouitine, H.~M. Gomes, J.~Read, T.~Abdessalem, and A.~Bifet.
\newblock River: machine learning for streaming data in python.
\newblock \emph{Journal of Machine Learning Research}, 22\penalty0
  (110):\penalty0 1--8, 2021.
\newblock URL \url{http://jmlr.org/papers/v22/20-1380.html}.

\bibitem[Moreno-Torres et~al.(2012)Moreno-Torres, Raeder, Alaiz-Rodr{\'\i}guez,
  Chawla, and Herrera]{moreno2012unifying}
J.~G. Moreno-Torres, T.~Raeder, R.~Alaiz-Rodr{\'\i}guez, N.~V. Chawla, and
  F.~Herrera.
\newblock A unifying view on dataset shift in classification.
\newblock \emph{Pattern recognition}, 45\penalty0 (1):\penalty0 521--530, 2012.

\bibitem[Nishida and Yamauchi(2007)]{nishida2007detecting}
K.~Nishida and K.~Yamauchi.
\newblock Detecting concept drift using statistical testing.
\newblock In \emph{International conference on discovery science}, pages
  264--269. Springer, 2007.

\bibitem[Nuyttens and Perrakis(2022)]{nannyml}
N.~Nuyttens and N.~Perrakis.
\newblock Nannyml.
\newblock \url{https://github.com/NannyML/nannyml}, 2022.
\newblock {Online; accessed 23-7-2023}.

\bibitem[Page(1954)]{page1954continuous}
E.~S. Page.
\newblock Continuous inspection schemes.
\newblock \emph{Biometrika}, 41\penalty0 (1/2):\penalty0 100--115, 1954.

\bibitem[Paszke et~al.(2019)Paszke, Gross, Massa, Lerer, Bradbury, Chanan,
  Killeen, Lin, Gimelshein, Antiga, et~al.]{paszke2019pytorch}
A.~Paszke, S.~Gross, F.~Massa, A.~Lerer, J.~Bradbury, G.~Chanan, T.~Killeen,
  Z.~Lin, N.~Gimelshein, L.~Antiga, et~al.
\newblock Pytorch: An imperative style, high-performance deep learning library.
\newblock \emph{Advances in neural information processing systems}, 32, 2019.

\bibitem[Pears et~al.(2014)Pears, Sakthithasan, and Koh]{pears2014detecting}
R.~Pears, S.~Sakthithasan, and Y.~S. Koh.
\newblock Detecting concept change in dynamic data streams.
\newblock \emph{Machine Learning}, 97\penalty0 (3):\penalty0 259--293, 2014.

\bibitem[Pearson(1900)]{pearson1900x}
K.~Pearson.
\newblock X. on the criterion that a given system of deviations from the
  probable in the case of a correlated system of variables is such that it can
  be reasonably supposed to have arisen from random sampling.
\newblock \emph{The London, Edinburgh, and Dublin Philosophical Magazine and
  Journal of Science}, 50\penalty0 (302):\penalty0 157--175, 1900.

\bibitem[Qahtan et~al.(2015)Qahtan, Alharbi, Wang, and Zhang]{qahtan2015pca}
A.~A. Qahtan, B.~Alharbi, S.~Wang, and X.~Zhang.
\newblock A pca-based change detection framework for multidimensional data
  streams: Change detection in multidimensional data streams.
\newblock In \emph{Proceedings of the 21th ACM SIGKDD International Conference
  on Knowledge Discovery and Data Mining}, pages 935--944, 2015.

\bibitem[Raab et~al.(2020)Raab, Heusinger, and Schleif]{raab2020reactive}
C.~Raab, M.~Heusinger, and F.-M. Schleif.
\newblock Reactive soft prototype computing for concept drift streams.
\newblock \emph{Neurocomputing}, 416:\penalty0 340--351, 2020.

\bibitem[Rabanser et~al.(2019)Rabanser, G{\"u}nnemann, and
  Lipton]{rabanser2019failing}
S.~Rabanser, S.~G{\"u}nnemann, and Z.~Lipton.
\newblock Failing loudly: An empirical study of methods for detecting dataset
  shift.
\newblock \emph{Advances in Neural Information Processing Systems}, 32, 2019.

\bibitem[Roberts(1959)]{10.2307/1266443}
S.~W. Roberts.
\newblock Control chart tests based on geometric moving averages.
\newblock \emph{Technometrics}, 1\penalty0 (3):\penalty0 239--250, 1959.
\newblock ISSN 00401706.
\newblock URL \url{http://www.jstor.org/stable/1266443}.

\bibitem[Ross et~al.(2012)Ross, Adams, Tasoulis, and
  Hand]{ross2012exponentially}
G.~J. Ross, N.~M. Adams, D.~K. Tasoulis, and D.~J. Hand.
\newblock Exponentially weighted moving average charts for detecting concept
  drift.
\newblock \emph{Pattern recognition letters}, 33\penalty0 (2):\penalty0
  191--198, 2012.

\bibitem[Roux et~al.(2022)Roux, Martin, and Bouché]{eurybia}
N.~Roux, J.~Martin, and T.~Bouché.
\newblock Eurybia.
\newblock \url{https://github.com/MAIF/eurybia}, 2022.
\newblock {Online; accessed 23-7-2023}.

\bibitem[Rubner et~al.(2000)Rubner, Tomasi, and Guibas]{rubner2000earth}
Y.~Rubner, C.~Tomasi, and L.~J. Guibas.
\newblock The earth mover's distance as a metric for image retrieval.
\newblock \emph{International journal of computer vision}, 40\penalty0
  (2):\penalty0 99--121, 2000.

\bibitem[Sakthithasan et~al.(2013)Sakthithasan, Pears, and
  Koh]{sakthithasan2013one}
S.~Sakthithasan, R.~Pears, and Y.~S. Koh.
\newblock One pass concept change detection for data streams.
\newblock In \emph{Pacific-Asia conference on knowledge discovery and data
  mining}, pages 461--472. Springer, 2013.

\bibitem[Scholz and Stephens(1987)]{scholz1987k}
F.~W. Scholz and M.~A. Stephens.
\newblock K-sample anderson--darling tests.
\newblock \emph{Journal of the American Statistical Association}, 82\penalty0
  (399):\penalty0 918--924, 1987.

\bibitem[Sethi and Kantardzic(2017)]{SETHI201777}
T.~S. Sethi and M.~Kantardzic.
\newblock On the reliable detection of concept drift from streaming unlabeled
  data.
\newblock \emph{Expert Systems with Applications}, 82:\penalty0 77--99, 2017.
\newblock ISSN 0957-4174.
\newblock \doi{https://doi.org/10.1016/j.eswa.2017.04.008}.
\newblock URL
  \url{https://www.sciencedirect.com/science/article/pii/S0957417417302439}.

\bibitem[Storkey(2009)]{storkey2009training}
A.~J. Storkey.
\newblock When training and test sets are different: characterizing learning
  transfer.
\newblock \emph{Dataset shift in machine learning}, 30\penalty0
  (3-28):\penalty0 6, 2009.

\bibitem[Swain and Ballard(1991)]{swain1991color}
M.~J. Swain and D.~H. Ballard.
\newblock Color indexing.
\newblock \emph{International journal of computer vision}, 7\penalty0
  (1):\penalty0 11--32, 1991.

\bibitem[Upton(1992)]{upton1992fisher}
G.~J. Upton.
\newblock Fisher's exact test.
\newblock \emph{Journal of the Royal Statistical Society: Series A (Statistics
  in Society)}, 155\penalty0 (3):\penalty0 395--402, 1992.

\bibitem[Van~Looveren et~al.(2019)Van~Looveren, Klaise, Vacanti, Cobb,
  Scillitoe, Samoilescu, and Athorne]{alibi-detect}
A.~Van~Looveren, J.~Klaise, G.~Vacanti, O.~Cobb, A.~Scillitoe, R.~Samoilescu,
  and A.~Athorne.
\newblock Alibi detect: Algorithms for outlier, adversarial and drift
  detection, 2019.
\newblock URL \url{https://github.com/SeldonIO/alibi-detect}.

\bibitem[Van~Rossum et~al.(2001)Van~Rossum, Warsaw, and Coghlan]{van2001pep}
G.~Van~Rossum, B.~Warsaw, and N.~Coghlan.
\newblock Pep 8--style guide for python code.
\newblock \emph{Python. org}, 1565:\penalty0 28, 2001.

\bibitem[Viehmann(2021)]{viehmann2021partial}
T.~Viehmann.
\newblock Partial wasserstein and maximum mean discrepancy distances for
  bridging the gap between outlier detection and drift detection.
\newblock \emph{arXiv preprint arXiv:2106.12893}, 2021.

\bibitem[Viehmann et~al.(2021)Viehmann, Antiga, Cortinovis, and
  Lozza]{torchdrift}
T.~Viehmann, L.~Antiga, D.~Cortinovis, and L.~Lozza.
\newblock Torchdrift.
\newblock \url{https://github.com/torchdrift/torchdrift}, 2021.
\newblock {Online; accessed 23-7-2023}.

\bibitem[Wang and Abraham(2015)]{wang2015concept}
H.~Wang and Z.~Abraham.
\newblock Concept drift detection for streaming data.
\newblock In \emph{2015 international joint conference on neural networks
  (IJCNN)}, pages 1--9. IEEE, 2015.

\bibitem[Welch(1947)]{welch1947generalization}
B.~L. Welch.
\newblock The generalization of ‘student's’problem when several different
  population varlances are involved.
\newblock \emph{Biometrika}, 34\penalty0 (1-2):\penalty0 28--35, 1947.

\bibitem[Wu and Olson(2010)]{wu2010enterprise}
D.~Wu and D.~L. Olson.
\newblock Enterprise risk management: coping with model risk in a large bank.
\newblock \emph{Journal of the Operational Research Society}, 61\penalty0
  (2):\penalty0 179--190, 2010.

\bibitem[{\v{Z}}liobait{\.e} et~al.(2016){\v{Z}}liobait{\.e}, Pechenizkiy, and
  Gama]{vzliobaite2016overview}
I.~{\v{Z}}liobait{\.e}, M.~Pechenizkiy, and J.~Gama.
\newblock An overview of concept drift applications.
\newblock \emph{Big data analysis: new algorithms for a new society}, pages
  91--114, 2016.

\end{thebibliography}

\documentclass[twoside,11pt]{article}

\usepackage[abbrvbib, preprint]{jmlr2e}

\usepackage[inline]{enumitem}
\usepackage{multirow}

\ShortHeadings{Frouros: A Python library for drift detection in machine learning systems}{Céspedes Sisniega and López García}
\firstpageno{1}

\begin{document}

\title{Frouros: A Python library for drift detection in machine learning systems}

\author{\name Jaime Céspedes Sisniega \email cespedes@ifca.unican.es \\
       \name Álvaro López García \email aloga@ifca.unican.es \\
       \addr Instituto de Física de Cantabria (IFCA), CSIC-UC, Spain
}
\editor{}

\maketitle

\begin{abstract}
    \verb+Frouros+ is an open-source Python library capable of detecting drift in machine learning systems. It provides a combination of classical and more recent algorithms for drift detection: both concept and data drift. We have designed it with the objective of making it compatible with any machine learning framework and easily adaptable to real-world use cases. The library is developed following a set of best development and continuous integration practices to ensure ease of maintenance and extensibility. The source code is available at \url{https://github.com/IFCA/frouros}.
\end{abstract}

\begin{keywords}
  Machine learning, Drift detection, Concept drift, Data drift, Python
\end{keywords}

\section{Introduction}

When deploying machine learning models in real-world applications, there is often the erroneous assumption that a given model will be used in a stationary environment, assuming that the same concepts learned during the training phase will remain valid at inference time \citep{gama2004learning}, or that training samples and production-time samples will come from the same distribution \citep{ackerman2021automatically}. However, in real-world scenarios, this is often far from being always true, and both situations may result in some type of \textit{drift} that can end up affecting the model performance \citep{vzliobaite2016overview}. Additionally, due to the high cost of collecting and labeling samples, this performance loss can often not be confirmed in many real-world problems, and other methods that only rely on distribution changes must be used.

In this paper, we present \verb+Frouros+, an open-source Python library for drift detection in machine learning systems. The library tries to fulfill two main objectives:
\begin{enumerate*} 
    \item to easily integrate into a machine learning system workflow that uses any machine learning framework, thus making it framework-agnostic;
    \item to unify in a single library the \textit{concept drift} detection part (traditionally researched and used for streaming/evolving data streams and incremental learning as described by \citealp{khamassi2018discussion}) with the research of change detection in the covariate distributions (also known as \textit{dataset shift} or \textit{distribution shift}, related to the field of statistical two-sample testing as introduced in \citealp{rabanser2019failing}, and methods that measure the distance between distributions as described by \citealp{goldenberg2019survey}).
\end{enumerate*}

\section{Drift detection}
\label{sec:drift-detection}

Traditionally there has been little consensus on the terminology and definitions of the different types of \textit{drift}, as stated by \cite{moreno2012unifying}. In order to adopt some clear definitions for the remainder of this paper, we apply those used by \cite{gama2014survey} for the \textit{concept drift} part, in combination with those used by \cite{rabanser2019failing} work for detecting \textit{dataset shift} using only the covariates. Therefore, we set up the following definitions assuming two different time points, ${t}$ and ${t + w}$, where $t$ could be any point in time, and $w$ could be the time at which the existence of change is checked \citep{bayram2022concept}. Thus, given $P(X, y) = P(y|X)P(X)$, a change in the joint distribution between two different times that can result in performance degradation can be described as $P_{t}(X, y) \neq P_{t + w}(X, y)$. The types of changes that can lead to degradation of the model's performance are categorized as follows:

\paragraph{Concept drift.} There is a change in the conditional probability $P(y|X)$, with or without a change in $P(X)$. Thus, it can be defined as $P_{t}(y|X) \neq P_{t + w}(y|X)$. Also known as \textit{real concept drift} \citep{gama2014survey}.

\paragraph{Data drift.} There is a change in $P(X)$. Therefore, this type of drift only focuses in the distribution of the covariates, $P(X)$, so $P_{t}(X) \neq P_{t + w}(X)$. Hereafter, we rename \textit{dataset shift} \citep{rabanser2019failing} to \textit{data drift}, in order to maintain consistency with the above definitions and with some of the related software mentioned in Section \ref{sec:comparison} that also refer to it as \textit{data drift} \citep{alibi-detect, menelaus}. A variation of \textit{data drift} can be considered \textit{label drift}, also known as \textit{prior probability shift} \citep{storkey2009training}, where the change occurs in the label distribution $P(Y)$ instead of the covariates $P(X)$. Thus, $P_{t}(Y) \neq P_{t + w}(Y)$.

\section{Overview and design}
\label{sec:overview-design}

The design and implementation of the library have been carried out with the aim of making it compatible with any machine learning framework, thus making it framework-agnostic. Detection methods are divided into the \textit{concept drift} and \textit{data drift} categories, depending on the type of drift they can detect, according to the definitions given in Section~\ref{sec:drift-detection}, and how they detect it. In terms of implementation, both categories share an abstract class named \verb+BaseDetector+ that contains the basic methods (reset) and attributes (callbacks) that a detectors must implement. \verb+Frouros+ also implements the well-known concept of \textit{callbacks}, presented in some machine learning frameworks, such as \verb+Keras+ \citep{chollet2015keras} and \verb+PyTorch Lightning+ \citep{Falcon_PyTorch_Lightning_2019}. This allows users to execute custom code at certain points. Additionally, we provide prequential error metrics \citep{gama2013evaluating} to evaluate the performance of \textit{concept drift} methods described in Section~\ref{subsec:concept-drift}.

\subsection{Concept drift}
\label{subsec:concept-drift}

These methods are aimed at detecting \textit{concept drift}. In order to update the detector, depending on the type of algorithm, they may require either the ground-truth labels/values or the values of the variable to be tracked, which is usually a model performance metric. The detector can receive the update values on demand; therefore, it is not mandatory to feed it with all the ground-truth labels/values or model performance metric values. All implemented methods work in a streaming manner, receiving one sample at a time.

In terms of implementation, each detector inherits from the \verb+BaseConceptDrift+ abstract class. Depending on the type of \textit{concept drift} detector, it also inherits from its corresponding type abstract class. The \verb+update+ method is used to update the detector's inner state, and it checks if drift is occurring each time a new value is received. Additionally, upon instantiation, the detector can optionally receive a list of \textit{callbacks}. It also receives a configuration class that contains specific parameters determining its behavior.

\subsection{Data drift}
\label{subsec:data-drift}

\textit{Data drift} methods focus on detecting changes by considering only the covariates. Therefore, these algorithms try to detect changes at a feature level by comparing new data distributions against reference data distributions.

Depending on the number of samples given to the detector to compare with the reference distribution, these methods can be classified as batch or streaming. The former uses a batch of samples to test against the reference distribution, and the latter can receive one sample at a time to update the inner state of the detector, therefore the samples must follow a sequential order.

Unlike \textit{concept drift} detectors, there is a \verb+fit+ method that stores the samples from the reference distribution used for comparison at inference time.
In the case of batch algorithms, the \verb+compare+ method is used to carry out the comparison and obtain a statistic or distance value depending on the type of detector. Whereas for the streaming detectors, an \verb+update+ method is used each time a new sample is added to the detector.

Based on the type of \textit{data drift} detector (statistical test or distance-based), checking if drift is taking place is done in different ways. After calling the \verb+compare+ or \verb+update+ methods, statistical tests provide the corresponding p-value that determines if drift exists. For distance-based detectors, the output is only the distance between the provided data and the data from the reference distribution. Therefore through the use of \textit{callbacks}, explained in Section \ref{subsec:callbacks}, p-values can be obtained. 

Moreover, \textit{data drift} detectors are implemented considering the type of data that they are expected to work with (categorical or numerical) and the number of features that can be considered (univariate or multivariate).

As mentioned in Section \ref{sec:drift-detection}, \textit{data drift} methods can also be used for detecting \textit{label drift} by considering the label distribution. Consequently, the described way of using the detectors for \textit{data drift} applies to \textit{label drift}.

\subsection{Callbacks}
\label{subsec:callbacks}

Execution of custom code at certain points is possible through the use of \textit{callbacks}. This adds enough flexibility to \verb+Frouros+ to be adapted to real-world use cases required by the users. \textit{Callbacks} are divided according to whether the detector belongs to the batch or streaming category.

For batch detectors, \textit{callbacks} inherit from \verb+BaseCallbackBatch+ class, which provides methods such as \verb+on_compare_start+ and \verb+on_compare_end+. These methods can be implemented to execute custom code before and after the comparison against the reference distribution is made.

In the case of streaming detectors, similar methods are provided for the \verb+update+ method. All \textit{callbacks} detectors share \verb+on_fit_start+ and \verb+on_fit_end+ provided by \verb+BaseCallback+, due to the fact that \textit{data drift} detectors can be categorised as batch or streaming.

Some examples of \textit{callbacks} already implemented and ready to use in the library include tracking the history of a \textit{concept drift} detector, computing p-values through the use of permutation tests in distance-based \textit{data drift} detectors, among others.

\section{Development}
\label{sec:development}

With the intention of following a set of open-source software development standards that allow the maintainability and extensibility of the library over time, we emphasize the following areas:

\paragraph{Continuous integration.} A continuous integration workflow based on \verb+GitHub+ \verb+Actions+ ensures that new modifications easily integrate with the existing code base and that they are compatible with multiple Python versions.

\paragraph{Documentation.} An API documentation is provided using \verb+sphinx+ and hosted on the \verb+Read+ \verb+the+ \verb+Docs+\footnote{\url{https://frouros.readthedocs.io}} website. The documentation includes some basic and advanced examples on the use of these detection methods.

\paragraph{Quality code.} To ensure minimum standards in terms of code quality, code coverage is set to be greater than 90\%, and some Python quality and style tools that comply with PEP8 standards \citep{van2001pep} are used, such as \verb+flake8+, \verb+pylint+, \verb+black+, and \verb+mypy+. 

\paragraph{Open-source.} In addition to the source code being available on \verb+GitHub+\footnote{\url{https://github.com/IFCA/frouros}}, \verb+Frouros+ package can be installed through the Python Package Index (PyPI)\footnote{\url{https://pypi.org/project/frouros}}. In terms of licensing, it is distributed under the BSD-3-Clause license.

\section{Comparison to related software}
\label{sec:comparison}

With regard to the \textit{concept drift} detection part, \verb+MOA+ \citep{bifet2010moa} has most of the methods that we are including, but they are implemented in Java. In Python, \verb+River+ \citep{JMLR:v22:20-1380}, which is focused on online machine learning and streaming data, offers some \textit{concept drift} detection methods, but only a subset of those presented here. Another Python library that contains this type of detectors is \verb+scikit-multiflow+ \citep{JMLR:v19:18-251}, but it has not been taken into account due to the fact that it was merged with the online machine learning library \verb+Creme+ \citep{creme}, resulting in the aforementioned \verb+River+ library.

For the \textit{data drift} section, \verb+Alibi Detect+ \citep{alibi-detect} has several algorithms related to the field of statistical two-sample hypothesis testing, and some of them can act both online (single sample) and offline (batch sample). \verb+TorchDrift+ \cite{torchdrift} also implements some statistical two-sample hypothesis testing methods, but in this case, it uses \verb+PyTorch+ \citep{paszke2019pytorch} for their implementation.

To the best of our knowledge, \verb+Menelaus+ \cite{menelaus} is the only open-source library that has both \textit{concept} and \textit{data drift} methods, although they classify them into the following types: \textit{change detection}, \textit{concept drift}, and \textit{data drift}. \textit{Concept drift} methods are implemented in such a way that the user must necessarily be in charge of controlling each iteration of the sample without offering some helper functions or classes to interact with the detector, as \verb+Frouros+ does with the \textit{callbacks} concept that is explained in Section~\ref{subsec:callbacks}.

Table \ref{tab:frameworks-algorithms} provides a more detailed view of the methods implemented in each of the libraries mentioned above, as well as those included in \verb+Frouros+. At the time of writing this paper, \verb+Frouros+ is listed as the library with the highest number of available methods with 28.

Moreover, several other libraries and tools have been excluded from Table \ref{tab:frameworks-algorithms} due to the fact that they implement a more limited number of methods or are more focused on building graphical dashboards and visual representations, such as \verb+Deepchecks+ \citep{deepchecks}, \verb+Eurybia+ \citep{eurybia}, \verb+Evidently+ \citep{evidently}, \verb+NannyML+ \citep{nannyml}, or \verb+UpTrain+ \citep{uptrain}.

\begin{table}[htbp]
\centering
\resizebox{0.84\linewidth}{!} {
\begin{tabular}{c|c|cccccc|c|}
\cline{2-9}
 & \textbf{Method} & {\rotatebox{90}{\textbf{Alibi-detect}}} & \rotatebox{90}{\textbf{Menelaus}} & \rotatebox{90}{\textbf{MOA}} & \rotatebox{90}{\textbf{River}} & \rotatebox{90}{\textbf{TorchDrift}} & \rotatebox{90}{\textbf{Frouros}} & \textbf{Reference} \\
\hline
\multicolumn{1}{|c|}{\multirow{18}{*}{\rotatebox[origin=c]{90}{\textbf{Concept drift}}}} & ADWIN & & \checkmark & \checkmark & \checkmark & & \checkmark & \cite{bifet2007learning} \\
\multicolumn{1}{|c|}{} & BOCD & & & & & & \checkmark & \cite{adams2007bayesian} \\
\multicolumn{1}{|c|}{} & CUSUM & & \checkmark & \checkmark & & & \checkmark & \cite{page1954continuous} \\
\multicolumn{1}{|c|}{} & DDM & & \checkmark & \checkmark & \checkmark & & \checkmark & \cite{gama2004learning} \\
\multicolumn{1}{|c|}{} & ECDD-WT & & & \checkmark & & & \checkmark & \cite{ross2012exponentially} \\
\multicolumn{1}{|c|}{} & EDDM & & \checkmark & \checkmark & \checkmark & & \checkmark & \cite{baena2006early} \\
\multicolumn{1}{|c|}{} & GMA & & & \checkmark & & & \checkmark & \cite{10.2307/1266443} \\
\multicolumn{1}{|c|}{} & HDDM-A & & & \checkmark & \checkmark & & \checkmark & \cite{frias2014online} \\
\multicolumn{1}{|c|}{} & HDDM-W & & & \checkmark & \checkmark & & \checkmark & \cite{frias2014online} \\
\multicolumn{1}{|c|}{} & KSWIN & & & & \checkmark & & \checkmark & \cite{raab2020reactive} \\
\multicolumn{1}{|c|}{} & LFR & & \checkmark & & & & & \cite{wang2015concept} \\
\multicolumn{1}{|c|}{} & Page Hinkley & & \checkmark & \checkmark & \checkmark & & \checkmark & \cite{page1954continuous} \\
\multicolumn{1}{|c|}{} & RDDM & & & \checkmark & & & \checkmark & \cite{barros2017rddm} \\
\multicolumn{1}{|c|}{} & SEED & & & \checkmark & & & & \cite{huang2014detecting} \\
\multicolumn{1}{|c|}{} & SeqDrift1 & & & \checkmark & & & & \cite{sakthithasan2013one} \\
\multicolumn{1}{|c|}{} & SeqDrift2 & & & \checkmark & & & & \cite{pears2014detecting} \\
\multicolumn{1}{|c|}{} & STEPD & & \checkmark & \checkmark & & & \checkmark & \cite{nishida2007detecting} \\
\multicolumn{1}{|c|}{} & MD3 & & \checkmark & & & & & \cite{SETHI201777} \\
\hline
\multicolumn{1}{|c|}{\multirow{25}{*}{\rotatebox[origin=c]{90}{\textbf{Data drift}}} } & Anderson-Darling test & & & & & & \checkmark & \cite{scholz1987k} \\
\multicolumn{1}{|c|}{} & Bhattacharyya Distance & & & & & & \checkmark & \cite{bhattacharyya1946measure} \\
\multicolumn{1}{|c|}{} & C2ST & \checkmark & & & & & & \cite{lopez2016revisiting} \\
\multicolumn{1}{|c|}{} & CDBD & & \checkmark & & & & & \cite{lindstrom2013drift} \\
\multicolumn{1}{|c|}{} & Context-aware MMD & \checkmark & & & & & & \cite{cobb2022context} \\
\multicolumn{1}{|c|}{} & Cramér-von Mises test  & \checkmark & & & & & \checkmark & \cite{cramer1928composition} \\
\multicolumn{1}{|c|}{} & Earth Mover's Distance & & & & & \checkmark & \checkmark & \cite{rubner2000earth} \\
\multicolumn{1}{|c|}{} & Fisher’s Exact test & \checkmark & & & & & & \cite{upton1992fisher} \\
\multicolumn{1}{|c|}{} & Hellinger Distance & & & & & & \checkmark & \cite{hellinger1909neue} \\
\multicolumn{1}{|c|}{} & HDDDM & & \checkmark & & & & & \cite{ditzler2011hellinger} \\
\multicolumn{1}{|c|}{} & Histogram Intersection & & & & & & \checkmark & \cite{swain1991color} \\
\multicolumn{1}{|c|}{} & Incremental KS test & & & & & & \checkmark & \cite{dos2016fast} \\
\multicolumn{1}{|c|}{} & JS Divergence & & & & & & \checkmark & \cite{lin1991divergence} \\
\multicolumn{1}{|c|}{} & kdq-Tree & & \checkmark & & & & & \cite{dasu2006information} \\
\multicolumn{1}{|c|}{} & KL Divergence & & & & & & \checkmark & \cite{kullback1951information} \\
\multicolumn{1}{|c|}{} & Kolmogorov-Smirnov test & \checkmark & & & & \checkmark & \checkmark & \cite{massey1951kolmogorov} \\
\multicolumn{1}{|c|}{} & Learned Kernel MMD & \checkmark & & & & & & \cite{liu2020learning} \\
\multicolumn{1}{|c|}{} & LSDD & \checkmark & & & & & & \cite{bu2016pdf} \\
\multicolumn{1}{|c|}{} & Mann-Whitney U test & & & & & & \checkmark & \cite{mann1947test} \\
\multicolumn{1}{|c|}{} & MMD & \checkmark & & & & \checkmark & \checkmark & \cite{JMLR:v13:gretton12a} \\
\multicolumn{1}{|c|}{} & NN-DVI & & \checkmark & & & & & \cite{liu2018accumulating} \\
\multicolumn{1}{|c|}{} & Partial MMD & & & & & \checkmark & & \cite{viehmann2021partial} \\
\multicolumn{1}{|c|}{} & PCA-CD & & \checkmark & & & & & \cite{qahtan2015pca} \\
\multicolumn{1}{|c|}{} & PSI & & & & & & \checkmark & \cite{wu2010enterprise} \\
\multicolumn{1}{|c|}{} & Welch's t-test & & & & & & \checkmark & \cite{welch1947generalization} \\
\multicolumn{1}{|c|}{} & ${\chi}^2$ & \checkmark & & & & & \checkmark & \cite{pearson1900x} \\
\hline
& {\textbf{\# Methods}} & 9 & 13 & 14 & 7 & 4 & \textbf{28} & \\
\cline{2-9}
\end{tabular}}
\caption{Drift detection methods by library}
\label{tab:frameworks-algorithms}
\end{table}

\section{Conclusion and future work}
\label{sec:conclusion}

This paper presents \verb+Frouros+, an open-source Python library for drift detection in machine learning systems that can be used with any machine learning framework, both for methods that aim to detect \textit{concept drift} and those that try to detect \textit{data drift}. Moreover, this library provides enough flexibility to adapt to the user's workflow through the so-called \textit{callbacks} concept, which is also used in other well-known machine learning libraries. Additionally, \verb+Frouros+ aims to meet open-source software development standards that allow for easy extension, both in terms of adding new methods or \textit{callbacks}.

For future work, in addition to regularly adding new detection methods, we plan to improve the performance of \textit{concept drift} methods using Python code optimizers like Cython \cite{behnel2010cython} and adapt \textit{data drift} methods to be used on GPUs. Furthermore, adding new \textit{callbacks}, as described in Section \ref{subsec:callbacks}, would enable the library to handle even more real-world use cases.


\acks{The authors acknowledge the funding from the Agencia Estatal de Investigación, Unidad de Excelencia María de Maeztu, ref. MDM-2017-0765 and the support from the project AI4EOSC “Artificial Intelligence for the European Open Science Cloud” that has
received funding from the European Union’s Horizon Europe Research and Innovation Programme under grant agreement 101058593.}


\newpage
\bibliography{bibliography}

\end{document}